The population mean and standard deviation are given below. Find the required probability and determine whether the given sample mean would be considered unusual. For a sample of n equals 65​, find the probability of a sample mean being greater than 227 if mu equals 226 and sigma equals 3.6. For a sample of n equals 65​, the probability of a sample mean being greater than 227 if mu equals 226 and sigma equals 3.6 is
Answer #1

Solution :

Given that ,

mean = \mu = 226

standard deviation = \sigma = 3.6

n = 65

\mu\bar x = 226

\sigma\bar x = \sigma / \sqrt n = 3.6 / \sqrt 65=0.45  

P(\bar x >227 ) = 1 - P(\bar x <227 )

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (227-226) / 0.45]

= 1 - P(z < 2.22)

Using z table,    

= 1 - 0.9868

probability = 0.0132